Shayna left home and traveled toward the ferry office at an average speed of 40 mph. Sometime later Mary left traveling in the same direction but at an average speed of 50 mph. After traveling for four hours Mary caught up with Shayna. Find the number of hours Shayna traveled before Mary caught up.
Answer by mananth(16946)

When Mary catches up both have traveled same distance from starting point
Let Shayna travel t hours before Mary starts
Shayna speed = 40 mph
distance = 40*(t+4)
Mary distance = 50 *40
40(t+4)=200
40t+160=200
40t=40
t=1
Shayna traveled 1 hour before Mary started